OpenStudy (anonymous):

An airline offers a discount one-way flight to New York City for $162.80. If the usual price for the ticket is $185.00, what is the percent discount? 12% 10% 15% 25%

OpenStudy (perl):

You can solve the equation 185 ( 1 - d/100) = 162,80

OpenStudy (anonymous):

discount = Usual Price - Amount Paid discount = 185 - 162.8 discount = $22.2 let Dp = percent discount \[Dp=\frac{ discount }{ UsualPrice }*100\] \[Dp=\frac{ 22.2 }{ 185 }*100\] Dp = 12% percent discount is 12%